OpenStudy (anonymous):

Can anyone teach me how to solve: 3^x=1 Without LOGS!!!

OpenStudy (anonymous):

1 can be written as 3^0 so 3^x = 3^0 on comparing x=0

OpenStudy (anonymous):

@sufiyancs is it right????

OpenStudy (anonymous):

Ohhh... Darn i forgot the basic rule... hehehe thanx @niksva

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Amm @niksva so how are we gonna solve 16^z=2

OpenStudy (anonymous):

write 16^z as 2^4z bcoz 2^4 = 16

OpenStudy (anonymous):

rest u know how to do it

OpenStudy (amistre64):

\[\frac{3^x}{3^x}=1\] \[3^{x-x}=1\] \[3^{0}=1\] \[x=0\]

OpenStudy (anonymous):

So u mean 4z=1 z=1/4 Right... Therefore, 16^(1/4) = 2 Thanxxx Again!